Crime data from the U.S. Federal Bureau of Investigation for the years 2012 and 2013 are given in the table below.

Year

Murder &       non-negligent manslaughter

Rape

Robbery

Aggravated assault

Burglary

Larceny-theft

Motor vehicle theft

Total

2012

14, 856

85, 141

355, 051

762, 009

2, 109, 932

6, 168, 874

723, 186

10, 219, 049

2013

14, 196

79, 770

345, 031

724, 149

1, 928, 465

6, 004, 453

699 ,594

9, 795 ,658

Total

29, 052

164, 911

700, 082

1, 486, 158

4 038, ,397

12, 173, ,327

1, 422 780

20, 014, 707

What is the probability that a randomly selected crime happened in 2013 or was rape
Answer #1

P(total crime in 2013) = 9, 795 ,658/20014707

P(total rape) =164, 911/20014707

P(total rape in 2013)=79, 770/20014707

P(crime happened in 2013 or was rape) = P(crime in 2013) + P(rape) - P(crime in 2013 and rape) = (9795658/20014707) +(164911/20014707 - 79770/20014707=0.4937